OpenStudy (anonymous):

which equation represents y= -x^2-10x-20 in vertex form?

OpenStudy (anonymous):

r u given some options or u wanna express it?

OpenStudy (anonymous):

Vertex equation is h=-b/2a Where h is (h,k) of vertex and a,b is ax^2+bx+c

OpenStudy (anonymous):

y= -)x-5)^2+5, y=-(x+5)^5+5, y=-_x-5)^2+15, y= -(x+50+10

OpenStudy (anonymous):

vertex is \[(-\frac{b}{2a},f(-\frac{b}{2a}))\]

OpenStudy (anonymous):

thats not a choice

OpenStudy (anonymous):

He wasn't attempting to answer it; he was giving you stuff to help you work up to the answer.

OpenStudy (anonymous):

oh so once you know the vertex, you can write it in that form easily without completing the square

OpenStudy (anonymous):

ok thank yall

OpenStudy (anonymous):

i can show you if you like, because it is easy

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we compute \[-\frac{b}{2a}\] in this case a = -1, b = -10 and so \[-\frac{b}{2a}=-5\]

OpenStudy (anonymous):

so you know it is going to look like \[y=-(x-(-5))^2 + number\] i.e. \[y=-(x+5)^2 + number\] the reason for the - sign out front is that you started with \[-x^2\]

OpenStudy (anonymous):

finding the number is easy. just replace x in the original expression by -5 and see what you get, since it is the second coordinate of the vertex. in this case you get \[y=-(-5)^2-10\times -5 -20=-25+50-20=25-20=5\] all these 5s are just a coincidence. your answer is therefore \[y=-(x+5)^2+5\]

OpenStudy (anonymous):

okay thank u so much..

OpenStudy (anonymous):

again the simple way is compute -b/2a and then plug it in for x. you will write y = (x+b/2a)^2 +whatever you got